KAGRA information

1) IFO got troubles in Y arm GR PDH signal glitches. So, the observation mode was down from around 10:00. Nakano-kun and remote Enomoto-kun fixed it by adjusting GR phase lock gain.

2) IFO lock itself was recovered around 19:00, however, the sensitivity had large fluctuation around 200 Hz a lot and the range varied btw 300kpc ~ 500 kpc. In addition to this, power up of IMC output power from 2.5 W to 5.2 W unlocked the IFO.

We temporally started IFO operation with this condition.

We found several problems in IFO situation and they were fixed as this klog report : http://klog.icrr.u-tokyo.ac.jp/osl/?r=14176

Then IFO was almost recovered with stable sensitivity. IMC output power was still set around 2.7W.

3) The remained CAL tasks were done in this early morning.

4) From the next relock timing, we will took alignment not by SAL but PRFPMI_LSC_LOCKED_RF state, then set IMC output power more up to 5.2W ( usual setting), then 7W.

shift summary








General information from KAGRA office


KAGRA/Observation/KAGRA-1/ObservationApril7toApril21/16th (last edited 2020-04-16 09:47:33 by miyoki)